OpenStudy (anonymous):

The number line shows the distance in meters of two jellyfish, A and B, from a predator located at point X: A horizontal number line extends from negative 3 to positive 3. The point A is at negative 1.5, the point 0 is labeled as X, and the point labeled B is at 0.5. Write an expression using subtraction to find the distance between the two jellyfish. PLEASE PLEASE PLEASE HELP

OpenStudy (anonymous):

@Abhisar @Brainybeauty @Cydney_Morgan @dan815 @eric_d @ganeshie8 @HaileyD @insa @just_one_last_goodbye @Kainui @lolabieber @MrNood @nincompoop @ophercule @Preetha @quickstudent @RedNeckOutLaw @sbc4242 @taylordanielle @xoxoimher @YanaSidlinskiy @ziqbal103

OpenStudy (anonymous):

 That's the explanation of paragraph 2, now, what is the relation between A and B and how can you calculate it?

OpenStudy (anonymous):

i have no idea and can you please just write an expression about this cause im running late on time and if you do i will give you a metal and become one of your fans so please cause im running late on time

OpenStudy (anonymous):

@Umangiasd

OpenStudy (anonymous):

@Abhisar @Brainybeauty @Cydney_Morgan @dan815 @eric_d @ganeshie8 @HaileyD @insa @just_one_last_goodbye @Kainui @lolabieber @MrNood @nincompoop @ophercule @Preetha @quickstudent @RedNeckOutLaw @sbc4242 @taylordanielle @xoxoimher @YanaSidlinskiy @ziqbal103

OpenStudy (anonymous):

You should learn to do this by yourself, it's pretty easy. As you have a real numbers line and the "spectator" is at 0, you can calculate the distance between A and B by simply substracting. distance=B-A=0.5m-(-1.5)m=2m